You are planning to buy a recondition F Premio car and for this purpose you have taken loan 1.5 million from Mutual

Trust Bank. Loan repayment time is 10 years. Assume 9% per year interest rate is applied. If you pay back the loan in ten

equal annual (per year) installments, how much will this installment be?


1
Expert's answer
2021-05-07T09:15:28-0400

Here, "P=1.5\times 10^6"


"r=9%, x" %, "t=10years"


Amount after 10 yeras"=\dfrac{p\times r\times t}{100}=\dfrac{1.5\times 10^6\times 9\times 10}{100}=1350000"


Installment for 1 year="\dfrac{1350000}{10}=135000=0.135 milion"